### v1

# Implementing a Robust Test of Galaxy Catalogue Completeness for Dark Siren Measurements of the Hubble Constant

Abstract
We present the application of a robust test of galaxy catalogue completeness to the gwcosmo pipeline. The method implements a straightforward statistical test for determining the apparent magnitude completeness limit of a magnitude-redshift sample. This offers an improved, less conservative approach compared with how galaxy catalogue completeness is currently estimated in the gwcosmo gravitational wave cosmology pipeline for determining the Hubble constant $H_{0}$. The test also does not require prior knowledge of the luminosity function, and thus returns a more robust estimate of the limiting apparent magnitude for a magnitude-redshift sample of galaxies. For GWTC-1 results using $B_{J}$-band photometry of galaxies in the GLADE catalogue, we find a $3.4\%$ improvement on the inference of $H_{0}$ using dark sirens only and a $1.3\%$ improvement for the combined posterior with GW170817. Using GLADE+, there is a $8.6\%$ improvement with dark sirens only and a $6.3\%$ improvement for the combined posterior with GW170817. However, the final posterior on $H_{0}$ using the GWTC-3 dataset with the GLADE+ $K$-band shows no improvement when applying the robust method. This is because the GLADE+ galaxy catalogue provides little or no coverage in the $K$-band for any of the GWTC-3 events. With the use of deeper galaxy catalogues in future gravitational wave cosmology analyses, the adoption of a less conservative estimate of magnitude completeness will become increasingly important.
keywords:
gravitational waves – cosmological parameters – methods: data analysis – catalogues
pubyear: 2025pagerange: Implementing a Robust Test of Galaxy Catalogue Completeness for Dark Siren Measurements of the Hubble ConstantImplementing a Robust Test of Galaxy Catalogue Completeness for Dark Siren Measurements of the Hubble Constant
1 Introduction
[…]
Over the past decade, gravitational-wave (GW) cosmology has emerged as a potentially powerful tool for resolving the current tension between early- and local- universe measurements of $H_{0}$.
One method for determining the Hubble constant from gravitational wave signals is the so-called "galaxy catalogue method", which was first proposed in Schutz (1986). Compact binary coalescences (CBCs) are self-calibrating distance indicators, yielding absolute distance measurements from analysis of their waveforms; they are referred to as standard sirens, the gravitational wave analogues to standard candles. Redshift information is degenerate with chirp mass in CBCs; therefore, their redshift $z$ must be obtained through other means. There are a number of methods for obtaining redshift information associated with dark sirens, several of which have been applied to real or simulated data to investigate the efficacy of bright and dark sirens for constraining the Hubble constant with current or future cosmological data (Chernoff & Finn, 1993; Taylor & Gair, 2012; MacLeod & Hogan, 2008; Messenger & Read, 2012; Nishizawa, 2017; Chen et al., 2018; Soares-Santos et al., 2019; Fishbach et al., 2019; Farr et al., 2019; Palmese et al., 2020; Mukherjee et al., 2021; Gray et al., 2023; Mukherjee et al., 2024; Afroz & Mukherjee, 2024). Here, we focus on the use of electromagnetic galaxy catalogues to obtain redshift information for GW events. Where a host galaxy cannot be identified through the detection of an electromagnetic counterpart to the GW event, one can statistically infer $H_{0}$ through the use of a galaxy catalogue. By assigning a probability to the potential host galaxies for each event, marginalising over these potential host galaxies, and combining the results for many events, a precise value for $H_{0}$ can be obtained (Schutz, 1986; Del Pozzo, 2012).
The success of the LIGO-Virgo-KAGRA (hereafter LVK) network of ground-based GW interferometers in detecting signals from compact binary mergers has made GW cosmology using CBCs as standard sirens a reality. In particular, the detection of the binary neutron star GW170817 and its electromagnetic counterpart provided our best individual constraint so far on $H_{0}$ from standard sirens, with an initial result of $H_{0}=70.0^{+12.0}_{-8.0}{\text{ km s}}^{-1}{\text{ Mpc}}^{-1}$ (Abbott et al., 2017b). To date, the best LVK constraints on $H_{0}$ from combining all detected standard sirens is $H_{0}=68^{+8}_{-6}{\text{ km s}}^{-1}{\text{ Mpc}}^{-1}$ (Abbott et al., 2023b).
While bright sirens – i.e. standard sirens with EM counterparts – remain our most powerful tool for inferring the value of $H_{0}$ from GW observations, only one such event has been detected to date. We must therefore make use of dark sirens to refine the posterior on $H_{0}$.
A full description of the galaxy catalogue method for inferring $H_{0}$ from dark sirens, as implemented in the python package gwcosmo version 1.0.0, can be found in Gray et al. (2020). It closely follows and extends the method outlined in Schutz (1986).
One crucial step in the aforementioned method is to determine the probability that the host galaxy of an event is contained within the galaxy catalogue used for analysis. This is dependent on, amongst other factors, the completeness of the galaxy catalogue. It is also important not to introduce selection effects in assigning probabilities to potential host galaxies. Because of this, observed galaxies that are fainter than the apparent magnitude threshold adopted are discarded in the analysis. This avoids introducing a bias towards brighter galaxies as potential hosts in regions of the catalogue where fainter galaxies are not observed due to the flux-limited nature of the surveys. [^1] Galaxy catalogue incompleteness can also impact adversely on gravitational-wave parameter estimation with imprecisely localised events. See e.g. Mo et al. (2024). A careful analysis to estimate the magnitude threshold for any sample of galaxies is therefore necessary.
GLADE and GLADE+ (Dálya et al., 2018, 2022) are composite catalogues made up of several surveys of varying depth and coverage. The GLADE+ catalogue comprises the GWGC, 2MPZ, 2MASS XSC, HyperLEDA and WISExSCOSPZ galaxy catalogues, and contains 22.5 million galaxies. Also included is the SDSS-DR16Q quasar catalogue. The GLADE+ catalogue is complete up to a luminosity distance of $d_{L}=47^{+4}_{-2}$ Mpc and contains bright galaxies up to 90% of the total expected $B$-band and $K$-band luminosities up to $\sim$130Mpc (Dálya et al., 2022). The previous GLADE catalogue does not contain the WISExSCOSPZ galaxy survey, and is complete to $d_{L}=37^{+3}_{-4}$ Mpc of the cumulative $B$-band galaxy luminosity. It contains $\sim$ 3 million objects that are categorised as galaxies (Dálya et al., 2018).
Determining the completeness of a galaxy catalogue is not straightforward. In the context of GW cosmology catalogue incompleteness can be treated in a number of ways (Palmese et al., 2023; Dalang & Baker, 2024). In this paper we define the completeness of a galaxy catalogue as an inherent characteristic of a flux-limited survey, which can be modelled through the identification of a limiting apparent magnitude to which the catalogue is considered complete. The flux-limited nature of a galaxy catalogue will mean that fainter galaxies are only observable nearby, while brighter galaxies will be missing from the survey at increasing distances.
The current implementation of gwcosmo uses the median apparent magnitude of the galaxy sample to define the limiting apparent magnitude in the galaxy catalogue method. While this conservative estimate seeks to avoid any biases due to selection effects, it also discards information about the potential host galaxies of dark sirens.
In this paper, we describe the implementation within gwcosmo 1.0.0 of a robust test for determining the limiting apparent magnitude of a galaxy catalogue. The test is applied to the pixelated version of the pipeline, as described in Gray et al. (2022). The pixelated implementation of gwcosmo introduced the treatment of galaxy catalogue completeness as directionally-dependent.
The remainder of this paper is organised as follows. Section 2 describes the robust test and its implementation into gwcosmo. Section 3 presents results of the new analysis using the GWTC-1 and GWTC-3 catalogues, and section 4 presents a discussion of the results and future work to be carried out.
2 Methods
2.1 The statistical test
We apply the statistical test for determining galaxy catalogue completeness first outlined in Rauzy (2001), hereafter R01. We will henceforth refer to this test as the robust method for estimating galaxy catalogue completeness. The robust method allows for the rigorous inference of the limiting apparent magnitude of a redshift-magnitude sample of galaxies.
We present an overview of the method here. A full derivation can be found in R01. The test is related to a statistical test derived in Efron & Petrosian (1992). By comparing two samples from the galaxy catalogue itself, rather than comparing a sample to the expected number of galaxies, this method assumes no specific model for the luminosity function, though it is assumed that the luminosity function is the same everywhere.
The method defines a statistic $T_{C}$ that tests different limiting apparent magnitudes. Essentially, the statistic tests for whether or not the catalogue is emptier than expected for a given ‘trial’ limiting apparent magnitude $m_{\rm lim}$.
For each sample of galaxies $\left\{(m_{i},z_{i})\right\}\rm$, we first define an array of test magnitude thresholds $m_{\rm lim}$, to which we will apply the statistical test. For each trial magnitude threshold $m_{\rm lim}$ we can define a corresponding absolute magnitude $M_{\rm lim}(\mu)$ such that:
$M_{\rm lim}(\mu)=m_{\rm lim}-\mu\text{.}$ (1)
This defines the faintest absolute magnitude, $M_{\rm lim}(\mu)$, at which a galaxy with distance modulus $\mu$ would be visible in the sample.
For each galaxy with $(M_{i},\mu_{i})$, and assuming a trial magnitude limit $m_{\rm lim}$, we can define a random variable $\zeta_{i}$, defined as:
$\zeta=\frac{F(M)}{(F(M_{\rm lim}(\mu))}$ (2)
where $F$ is the cumulative luminosity function. The variable $\zeta$ is distributed within the interval $[0,1]$ and independent of the sample.
The cumulative luminosity function $F$ in equation 2 can be approximated by counting the number of galaxies in two regions of the sample.
For a galaxy with absolute magnitude $M_{i}$ this random variable effectively compares the number of galaxies with $\mu<\mu_{i}$ and $M<M_{i}$ to the number of galaxies with $\mu<\mu_{i}$ and up to the trial apparent magnitude limit $m_{\rm lim}$. The previous expression is therefore equivalent to:
$\zeta_{i}=\frac{r_{i}}{n_{i}+1}\text{,}$ (3)
where for each galaxy $i$ we have
$r_{i}=S_{1},\quad n_{i}=S_{1}\cup S_{2}\text{,}$ (4)
where $S_{1}$ is the area drawn out by $\mu<\mu_{i}$ and $M<M_{i}$, while $S_{2}$ is defined by $\mu>\mu_{i}$ and $M^{i}_{\rm lim}>M>M_{i}$. $r$ and $n$ are integrals corresponding to the number of galaxies in those areas. Figure 1 illustrates how the areas $S_{1}$ and $S_{2}$ are constructed, along with the $(M,\mu)$ samples, for a galaxy in one pixel of the $B$-band of GLADE+. In each subsample $S_{1}$ and $S_{2}$, $M$ and $\mu$ are, by construction, independent.
It is proven in R01 that the variables $\zeta_{i}$ are also independent of $\mu$ under the null hypothesis. The expectation $E_{i}$ and variance $V_{i}$ of each random statistic $\zeta_{i}$ are:
$E_{i}=\frac{1}{2}\text{ and }V_{i}=\frac{n_{i}-1}{12(n_{i}+1)}\text{.}$ (5)
R01 defines the quantity $T_{C}$ as:
$T_{C}=\frac{\sum\limits^{N}_{i=1}\Big{(}\zeta_{i}-\frac{1}{2}\Big{)}}{\sum\limits^{N}_{i=1}V_{i}}\text{,}$ (6)
Figure 1: Illustrating the $S_{1}$ and $S_{2}$ areas for a single pixel in the GLADE $B$-band. $M$ vs $\mu$ with $S_{1}$ and $S_{2}$ for a galaxy ($M_{i}$,$\mu_{i}$). The dash-dotted orange and blue lines show, respectively, the robust and median apparent magnitude thresholds for this sample of galaxies. The green line shows a test limit magnitude $m_{*}$ that is fainter than the true $m_{thr}$.
R01 goes on to show that the expectation of $T_{C}$ is zero and its variance is unity. $T_{C}$ can be estimated without any a-priori assumptions about the form of the luminosity function. If the sample is complete to the trial value $m_{\rm lim}$, then $T_{C}$ follows a Gaussian distribution around zero; $T_{C}$ starts going systematically negative when there is a deficit of galaxies fainter than $M_{\rm lim}(\mu)$. This deficit is caused by the trial magnitude $m_{\rm lim}$ being larger than the true magnitude threshold $m_{\rm thr}$.
The statistic $T_{C}$ becoming systematically very negative is therefore indicative of catalogue incompleteness. This is a result of the region $S_{2}$ being systematically emptier than the region $S_{1}$, due to the impact of the apparent magnitude limit. Following Rauzy (2001), we take $T_{C}=-3$ as a threshold value that indicates the limiting magnitude for catalogue completeness.
2.2 Implementation
The method is implemented into the pixelated version of gwcosmo, replacing the median apparent magnitude in each pixel as an estimate of the limiting magnitude for that pixel. A full overview of the pixelated method is available in Gray et al. (2022). In this method, galaxy catalogues are broken into directional pixels of equal area. Each "pixel" of the galaxy catalogue gives a magnitude-redshift sample of galaxies in a given range of right ascension and declination, for which we estimate a limiting apparent magnitude $m_{\rm thr}$.
The value of this magnitude threshold affects the resulting posterior on $H_{0}$ in two ways: firstly, all galaxies fainter than $m_{\rm thr}$ are discarded from the analysis. Secondly, it informs the probability of the host galaxy of the GW event being within the galaxy catalogue.
We summarise the most relevant parts of the equations that involve the probability of the host galaxy of the event being in the galaxy catalogue, and thus the apparent magnitude threshold $m_{\rm thr}$.
In order to obtain a final posterior on $H_{0}$, we need to marginalise the probability of the data $x_{GW}$ over the two propositions $G$ and $\overline{G}$, with:
  • g representing proposition $G$ and $\overline{G}$,
  • $G$ representing that the host galaxy is within the galaxy catalogue,
  • $\overline{G}$ representing that the host galaxy is outwith the galaxy catalogue.
In the gwcosmo pipeline, these depend on the apparent magnitude threshold of the catalogue. The term $\overline{G}$ integrates over the portion of the galaxy luminosity function that cannot be observed due to being fainter than the catalogue’s limiting apparent magnitude.
$p(x_{GW}|D,H_{0})=\sum_{g}p(x_{GW}|g,D,H_{0})p(g|D,H_{0})\text{.}$ (7)
$p(G|z,\Omega,M,m,D,s,H_{0})$ is approximated by a Heaviside step function around the apparent magnitude threshold $m_{\rm thr}$, corresponding to a sharp cut-off on apparent magnitude.
[…]
Equation 8 is, for a given Hubble constant $H_{0}$, the probability that the host galaxy for a gravitational wave event is contained within the catalogue. This is not the only part of the gwcosmo pipeline that is affected by the choice of $m_{\rm thr}$; galaxies with apparent magnitudes below the threshold are discarded in the analysis. This means that when approximating $m_{\rm thr}$ as the median apparent magnitude, half of the catalogue is discarded.
In the $B$-band, the GLADE and GLADE+ catalogues contain millions of galaxies, and each pixel contains thousands of galaxies. The complexity of the robust test described above goes up with $\mathcal{N}^{2}$, making it computationally expensive and slow to determine the apparent magnitude threshold of each pixel for each gravitational wave event. To overcome this, we randomly sample a subset of $\mathcal{N}=400$ galaxies with $(m_{i},z_{i})$ multiple times and evaluate $m_{\rm thr}$ for each subset of galaxies. The final value of the magnitude threshold $m_{\rm thr}$ for that pixel is then taken to be the mean of the evaluated thresholds for the different random sub-samples. This only needs to be done for the $B$-band of the GLADE and GLADE+ catalogues; in the $K$-band, each pixel is more sparsely populated.
Future instances of gwcosmo will compute all $m_{\rm thr}$ only once for entire catalogues rather than for each event, meaning that in the future this sub-sampling will not be necessary.
The redshifts $z_{i}$ of the galaxies in the sample have an associated photometric or spectroscopic uncertainty $\sigma_{z_{i}}$. To propagate redshift uncertainties to the magnitude threshold estimate, we make the following two assumptions:
  • The galaxies have a redshift
  • The redshift uncertainties are Gaussian
We therefore assume the uncertainties on the galaxy redshifts are described by a truncated Gaussian. The distribution that is sampled from for each redshift $z_{i}$ is truncated to only allow positive redshift values (we assume every galaxy has a redshift).
In order to sample over the redshift distributions, the magnitude threshold $m_{\rm thr}$ is estimated for several samples of $(m_{i},z_{i})$ for each pixel. This in turn returns a range of $m_{\rm thr}$ values. The apparent magnitudes $m_{i}$ have an associated uncertainty $\sigma_{m_{i}}$. However, for direct comparison to the median method of estimating incompleteness, which does not incorporate $m_{i}$ uncertainties, we choose to ignore them in this work.
The trial magnitude thresholds $m_{\rm lim}$ for each sample are taken to lie between $m_{\rm lim}=m_{\rm med}-0.5$ and $m_{\rm lim}=m_{\rm med}+4$ with a step of $0.05$ mag, where $m_{\rm med}$ is median apparent magnitude for each galaxy-redshift sample.
3 Results
The previously outlined robust method of estimating the apparent magnitude threshold $m_{thr}$ is implemented into gwcosmo and applied to the inference of $H_{0}$ using the GWTC-1 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b) catalogues of gravitational wave events. As was the case in the LVK analyses previously performed on those catalogues, only events with an SNR above $11$ are considered in this work.
3.1 GWTC-1
GWTC-1 is made up of all compact binary coalescences detected up to the end of the second observing run of the LVK network. (Abbott et al. (2019)) Of the eleven events in GWTC-1, seven are used in the inference of $H_{0}$. Six of them are binary black holes, and one of them is the BNS event GW170817 which had an associated EM counterpart; it is not affected by any changes to the catalogue method.
The GWTC-1 data has previously been used in the inference of $H_{0}$ using the GLADE catalogue. Here we repeat the analysis using the $B_{J}$-band of both the GLADE and GLADE+ catalogues. One caveat to note is that the assumption that the luminosity function is universal might fail when using the $B_{J}$-band of the GLADE+ catalogue; therefore, this analysis should serve as proof of principle for consequent analyses.
In the GWTC-1 analysis using the $B$-band of the GLADE and GLADE+ catalogues, there is a change in the recovered posterior on $H_{0}$ when using the robust method of estimating $m_{\rm thr}$ compared to the final posterior using $m_{\rm med}$ as the threshold. Figure 2 shows results for the GWTC-1 analysis using the GLADE $B$-band. The width of the final recovered posterior on $H_{0}$ using the robust method is 1.3% narrower than that of the median method for the 68.3% percentile when considering only dark sirens. The final posterior with GW170817 is 3.4% narrower using the robust method.
Figure 3 shows the same analysis using the GLADE+ $B$-band instead of the GLADE $B$-band. GW170814 remains unchanged, being analysed using the DES (Drlica-Wagner et al., 2018) catalogue. For this analysis, when using robust, the final posterior on $H_{0}$ is 8.6% narrower when considering only dark sirens, and 6.3% narrower when considering both dark and bright sirens. This is a clear improvement to the GWTC-1 results with GLADE+ when using the robust method.
Figure 4 shows the posterior on $H_{0}$ for the event GW170814. The $87$ deg2 sky localisation area of GW170814 was entirely contained within the DES footprint (Doctor et al., 2019). The DES-Y1 catalogue consists of $\sim$137 million objects over $\sim$1800 deg2 in the DES $grizY$ filters. The $10\sigma$ limiting magnitudes for galaxies are $g=23.4$, $r=23.2$, $i=22.5$, $z=21.8$ and $Y=20.1$ (Drlica-Wagner et al., 2018). The catalogue includes photometric redshift estimates. The gravitational wave event GW170814 originated from within the area mapped by the DES-Y1 survey. In the analysis of GW170814 with gwcosmo, we use the $g$-band data from DES, as it is the band in which the survey is most complete, with the $95\%$ completeness magnitude limit in the $g$- band quoted at $23.72$ mag in a sample of high quality objects (Abbott et al., 2018). Figure 5 shows posteriors on $H_{0}$ for each individual dark siren in the GWTC-1 catalogue.
3.2 GWTC-3
The GWTC-3 catalogue consists of 90 events, of which 47 are used in this analysis. Of these 47 events, 42 are BBHs, 2 are BNSs (the bright siren GW170817 and GW190425), 2 are NSBHs (GW200105 and GW200115) and one is the asymmetric mass binary GW190814 (Abbott et al., 2023a, 2017a, 2020a, 2021b, 2020b). In the original analysis presented in Abbott et al. (2023b), the $K$-band of the GLADE+ catalogue was found to be more appropriate than the $B_{J}$-band for analysis; it is less affected by galactic dust, and the behaviour of its luminosity function can therefore be better approximated. While this does not affect tests of galaxy catalogue completeness, it does affect the luminosity weighting of galaxies in the sample (Abbott et al., 2023b).
The completeness of the GLADE+ catalogue decreases more rapidly past $d_{L}\sim 100$Mpc in the $K$-band than in the $B$-band (Dálya et al., 2018). Applying the robust method to the GLADE+ catalogue with a pixel size of $N_{\rm side}=32$, the mean $K$-band apparent magnitude threshold is $m_{\rm thr}=13.49$. By comparison, the median method gives $m_{\rm thr}=12.91$. While the robust method allows us to use more galaxies, the apparent magnitude threshold is still comparatively bright, reducing the impact of the method on the recovered $H_{0}$ compared to when using the $B_{J}$-band.
Results from the GWTC-3 analysis are shown in figure 6. As anticipated, there is no improvement in the recovered posterior on $H_{0}$ from this analysis. Figure 7 shows results for individual event posteriors for both the robust and median methods. The recovered posteriors are similar for all events, due to the lack of coverage in the $K$-band at the luminosity distances at which potential host galaxies would be located.
Figure 2: Final posterior on $H_{0}$ using the GLADE $B_{J}$-band on the GWTC-1 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
Figure 3: Final posterior on $H_{0}$ using the GLADE+ $B_{J}$-band on the GWTC-1 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
Figure 4: Posterior on $H_{0}$ for the event GW170814 using the DES catalogue. The blue line shows results using the median apparent magnitude as $m_{\rm thr}$, and the orange line shows the final posterior when the robust method of inferring $m_{\rm thr}$ is used.
Figure 5: Results using the robust (blue) and median (orange) methods for individual events using the GLADE+ $B_{J}$-band with the GWTC-1 catalogue. The event GW170814 is analysed using the $g$-band of the DES-Y1 catalogue.
Figure 6: Final posterior on $H_{0}$ using the GLADE+ $K$-band on the GWTC-3 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
Figure 7: Inferred $H_{0}$ using the robust (orange) and median (blue) methods for estimating galaxy catalogue completeness, for individual events in the GWTC-3 catalogue. The $K$-band of the GLADE+ catalogue is used for analysis.
The final results for both GWTC-1 and GWTC-3 are summarised in table 1.
Catalogue Method $H_{0}$ (km s-1 Mpc-1) $H_{0}$ (km s-1 Mpc-1)
Dark Sirens with GW170817
                                GWTC-1, $B_{J}$-band
GLADE median $68.8^{+46.9}_{-21.7}$ $69.6^{+19.2}_{-8.3}$
GLADE robust $68.7^{+43.1}_{-24.6}$ $69.5^{+18.1}_{-8.5}$
GLADE+ median $65.7^{+41.8}_{-22.7}$ $69.3^{+17.2}_{-8.3}$
GLADE+ robust $67.9^{+35.1}_{-23.8}$ $69.4^{+16.1}_{-7.8}$
                                GWTC-3, $K$-band
GLADE+ median $66.8^{+12.9}_{-11.6}$ $68.6^{+8.4}_{-6.2}$
GLADE+ robust $67.7^{+13.7}_{-11.4}$ $68.9^{+8.5}_{-6.5}$
Table 1: Final results for constraints on $H_{0}$ in km s-1 Mpc-1 from the GWTC-1 and GWTC-3 datasets, using different methods and galaxy catalogues. The GWTC-1 dataset was analysed using the $B_{J}$-band of both the GLADE and GLADE+ catalogues, while the GWTC-3 dataset was analysed using the $K$-band of the GLADE+ catalogues. Confidence intervals are quoted at the $1\sigma$ level.
4 Conclusions
In this work we presented new results on the constraints on $H_{0}$ from dark sirens when we apply a robust test of completeness to the galaxy catalogue method.
There was no improvement to the posterior on $H_{0}$ with the robust method for the GWTC-3 analysis using the $K$-band of GLADE+. This is because the galaxy catalogue provides little or no coverage for any of the events in that band, whether the median or robust method is used. The final result is similar to the "empty catalogue" posteriors for each GWTC-3 event.
However, the final posterior on $H_{0}$ showed significant improvement when applying the robust method to the GWTC-1 analysis using the $B_{J}$-band of the GLADE+ catalogue. When only dark sirens were considered, the $1\sigma$ posterior was 8.6% narrower when using the robust method than when using the median apparent magnitude as a threshold. While the GLADE+ $B$-band is less reliable for tests of completeness due to the behaviour of its luminosity function [^2] See the discussion in (Abbott et al., 2023b), the result demonstrates the need for a careful treatment of the apparent magnitude threshold of future, deeper galaxy catalogues in order to obtain the best constraints on $H_{0}$ from dark standard sirens.
The robust method applied here is more computationally expensive than simply taking the median apparent magnitude as a threshold – with the complexity scaling as $\mathcal{N}^{2}$, where $\mathcal{N}$ is the number of galaxies in the sample. However, future instances of the pipeline will compute $m_{\rm thr}$ for the entire catalogue prior to analysis, circumventing the need to re-apply the method for each event. This will lead to improved performance and would eliminate the need for sub-sampling of galaxies. The threshold value used for determining $m_{\rm thr}$ from $T_{C}$ can also be refined in future work. Moreover, an uncertainty in the estimate of $m_{\rm thr}$ for each pixel can also be derived from the measurement uncertainties on apparent magnitudes and redshifts. These threshold uncertainties could then, in principle, also be incorporated into the gwcosmo pipeline. In this paper, the method was implemented into gwcosmo 1.0.0, but it will also be possible to incorporate it into version 2.0.0.
While the robust method does not require that we know the exact form of the luminosity function, it does still make the assumption that the luminosity function is universal for the galaxy catalogue band considered. This represents a caveat when applying the robust method to the $B$-band of the GLADE and GLADE+ galaxy catalogues, and further investigation of the validity of this assumption for other bands and other catalogues will be carried out in future work.
Ongoing work ahead of the fifth LVK observing run, O5, is exploring the quantitative effects of having a deeper apparent magnitude limit in galaxy catalogues used for gravitational wave cosmology. With deeper surveys, we predict that excessively conservative estimates for $m_{\rm thr}$ will have a greater impact, making the implementation of robust completeness methods increasingly important in future work. Mock data challenges with deeper EM galaxy catalogues ahead of O5 will allow us to quantify the effect of using robust for future analyses.
Our future work will focus on applying the robust method to mock data in order to fully characterise potential biases and explore the effect of a more rigorously and robustly defined $m_{\rm thr}$ on the inference of $H_{0}$ when analysing GW data with deeper galaxy catalogues. We will also extend analysis to other colour bands, including the $B_{J}$ band, seeking to exploit the property of the robust method that it does not require the adoption of a specific parametric form for the galaxy luminosity function. Moreover, our future work we will also extend the analysis presented here to the case of galaxy surveys described by both a faint and bright apparent magnitude limit, applying the robust completeness test first developed in Johnston et al. (2007).
Acknowledgements
The authors are grateful to En-Tzu Lin, Rachel Gray, Gavin Lamb, Surojit Saha, Surhud More, Maciej Bilicki, Gergely Dalya, Carl-Johan Haster and Suvodip Mukherjee for their helpful feedback and suggestions.
This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. L. D. was supported by the Science & Technology Facilities Council (ST/R504750/1). M. H. is supported by the Science and Technology Facilities Council (Ref. ST/L000946/1). We acknowledge the use of the following python packages in this work: gwcosmo (Gray et al., 2020), Matplotlib (Hunter, 2007), healpy (Górski et al., 2005; Zonca et al., 2019).
Data Availability
The GWTC-3 dataset is available from LIGO Scientific Collaboration and Virgo Collaboration and KAGRA Collaboration (2023). The GWTC-1 dataset is available from https://www.gw-openscience.org/GWTC-1. The GLADE and GLADE+ catalogues are available from the GLADE website http://glade.elte.hu/. The DES-Y1 catalogue is available from https://des.ncsa.illinois.edu/releases/y1a1.

### v2

@@
 The test also does not require prior knowledge of the luminosity function, and thus returns a more robust estimate of the limiting apparent magnitude for a magnitude-redshift sample of galaxies.
-For GWTC-1 results using $B_{J}$-band photometry of galaxies in the GLADE catalogue, we find a $3.4\%$ improvement on the inference of $H_{0}$ using dark sirens only and a $1.3\%$ improvement for the combined posterior with GW170817.
+For GWTC-1 results using $B$-band photometry of galaxies in the GLADE catalogue, we find a $3.4\%$ improvement on the inference of $H_{0}$ using dark sirens only and a $1.3\%$ improvement for the combined posterior with GW170817.
 Using GLADE+, there is a $8.6\%$ improvement with dark sirens only and a $6.3\%$ improvement for the combined posterior with GW170817.
@@
 The GWTC-1 data has previously been used in the inference of $H_{0}$ using the GLADE catalogue.
-Here we repeat the analysis using the $B_{J}$-band of both the GLADE and GLADE+ catalogues.
-One caveat to note is that the assumption that the luminosity function is universal might fail when using the $B_{J}$-band of the GLADE+ catalogue; therefore, this analysis should serve as proof of principle for consequent analyses.
+Here we repeat the analysis using the $B$-band of both the GLADE and GLADE+ catalogues.
+One caveat to note is that the assumption that the luminosity function is universal might fail when using the $B$-band of the GLADE+ catalogue; therefore, this analysis should serve as proof of principle for consequent analyses.
 In the GWTC-1 analysis using the $B$-band of the GLADE and GLADE+ catalogues, there is a change in the recovered posterior on $H_{0}$ when using the robust method of estimating $m_{\rm thr}$ compared to the final posterior using $m_{\rm med}$ as the threshold.
@@
 In the original analysis presented in Abbott et al.
-(2023b), the $K$-band of the GLADE+ catalogue was found to be more appropriate than the $B_{J}$-band for analysis; it is less affected by galactic dust, and the behaviour of its luminosity function can therefore be better approximated.
+(2023b), the $K$-band of the GLADE+ catalogue was found to be more appropriate than the $B$-band for analysis; it is less affected by galactic dust, and the behaviour of its luminosity function can therefore be better approximated.
 While this does not affect tests of galaxy catalogue completeness, it does affect the luminosity weighting of galaxies in the sample (Abbott et al., 2023b).
@@
 By comparison, the median method gives $m_{\rm thr}=12.91$.
-While the robust method allows us to use more galaxies, the apparent magnitude threshold is still comparatively bright, reducing the impact of the method on the recovered $H_{0}$ compared to when using the $B_{J}$-band.
+While the robust method allows us to use more galaxies, the apparent magnitude threshold is still comparatively bright, reducing the impact of the method on the recovered $H_{0}$ compared to when using the $B$-band.
 Results from the GWTC-3 analysis are shown in figure 6.
@@
 The recovered posteriors are similar for all events, due to the lack of coverage in the $K$-band at the luminosity distances at which potential host galaxies would be located.
-Figure 2: Final posterior on $H_{0}$ using the GLADE $B_{J}$-band on the GWTC-1 dataset.
+Figure 2: Final posterior on $H_{0}$ using the GLADE $B$-band on the GWTC-1 dataset.
 Vertical dashed lines show the $1\sigma$ intervals.
 The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
-Figure 3: Final posterior on $H_{0}$ using the GLADE+ $B_{J}$-band on the GWTC-1 dataset.
+Figure 3: Final posterior on $H_{0}$ using the GLADE+ $B$-band on the GWTC-1 dataset.
 Vertical dashed lines show the $1\sigma$ intervals.
@@
 The blue line shows results using the median apparent magnitude as $m_{\rm thr}$, and the orange line shows the final posterior when the robust method of inferring $m_{\rm thr}$ is used.
-Figure 5: Results using the robust (blue) and median (orange) methods for individual events using the GLADE+ $B_{J}$-band with the GWTC-1 catalogue.
+Figure 5: Results using the robust (blue) and median (orange) methods for individual events using the GLADE+ $B$-band with the GWTC-1 catalogue.
 The event GW170814 is analysed using the $g$-band of the DES-Y1 catalogue.
@@
 Dark Sirens with GW170817
-                                GWTC-1, $B_{J}$-band
+                                GWTC-1, $B$-band
 GLADE median $68.8^{+46.9}_{-21.7}$ $69.6^{+19.2}_{-8.3}$
@@
 Table 1: Final results for constraints on $H_{0}$ in km s-1 Mpc-1 from the GWTC-1 and GWTC-3 datasets, using different methods and galaxy catalogues.
-The GWTC-1 dataset was analysed using the $B_{J}$-band of both the GLADE and GLADE+ catalogues, while the GWTC-3 dataset was analysed using the $K$-band of the GLADE+ catalogues.
+The GWTC-1 dataset was analysed using the $B$-band of both the GLADE and GLADE+ catalogues, while the GWTC-3 dataset was analysed using the $K$-band of the GLADE+ catalogues.
 Confidence intervals are quoted at the $1\sigma$ level.
@@
 The final result is similar to the "empty catalogue" posteriors for each GWTC-3 event.
-However, the final posterior on $H_{0}$ showed significant improvement when applying the robust method to the GWTC-1 analysis using the $B_{J}$-band of the GLADE+ catalogue.
+However, the final posterior on $H_{0}$ showed significant improvement when applying the robust method to the GWTC-1 analysis using the $B$-band of the GLADE+ catalogue.
 When only dark sirens were considered, the $1\sigma$ posterior was 8.6% narrower when using the robust method than when using the median apparent magnitude as a threshold.
@@
 Our future work will focus on applying the robust method to mock data in order to fully characterise potential biases and explore the effect of a more rigorously and robustly defined $m_{\rm thr}$ on the inference of $H_{0}$ when analysing GW data with deeper galaxy catalogues.
-We will also extend analysis to other colour bands, including the $B_{J}$ band, seeking to exploit the property of the robust method that it does not require the adoption of a specific parametric form for the galaxy luminosity function.
+We will also extend analysis to other colour bands, including the $B$ band, seeking to exploit the property of the robust method that it does not require the adoption of a specific parametric form for the galaxy luminosity function.
 Moreover, our future work we will also extend the analysis presented here to the case of galaxy surveys described by both a faint and bright apparent magnitude limit, applying the robust completeness test first developed in Johnston et al.